OpenStudy (anonymous):

Can you please explain to me why any number to the power of 0 will always equal one? This 8th grade math is beating me

OpenStudy (anonymous):

did u mean : prove \[x^0=1\]

OpenStudy (anonymous):

I need to explain it to a 12 year old - I don't get it...

OpenStudy (anonymous):

actually i am also of the same age

OpenStudy (anonymous):

ok let me explain wait for a minute

OpenStudy (anonymous):

it's been a long time since I was in school! I've forgotten a lot

hero (hero):

You can only prove it using the rules of algebra. Kushashwa, just remember you're not average.

OpenStudy (anonymous):

i know that i am weak in maths hero ... i am trying to improve atelast upto average level

OpenStudy (anonymous):

this site has lead me to from weakest level to atleast weak level .. i am improving i think

hero (hero):

I meant that you're above average

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

ok, but I still don't understand why any number to the power of 0 will always = 1

OpenStudy (anonymous):

hmn i think this can be the proof see u know that 1 = a^n / a^n 1= a^(n-n) = a^0 a^0 = 1 hence a^0 = 1

hero (hero):

That won't mean anything to a twelve year old

OpenStudy (anonymous):

or a 38 y/o either lol

hero (hero):

It was a good attempt though

hero (hero):

You would have been better off using actual numbers

OpenStudy (anonymous):

yes a very good attempt...

OpenStudy (anonymous):

but it's not making sense to me

OpenStudy (anonymous):

1 * 4 * 4 * 4 = 4^3 1* 4 * 4 = 4^2 1 * 4 = 4^1 Therefore, 1 = 4^0

OpenStudy (anonymous):

Since... x^a*x^b=x^(a+b) --------------- laws of exponents hence x^a*x^0=x^(a+0)=x^a if x^a*x^0=x^a then, x^0 is a multiplicative identity, and must be equal to 1.

OpenStudy (anonymous):

this is the easiest one hero is this right ?

OpenStudy (anonymous):

@Tyler that is just a pattern not a proof i think that

hero (hero):

kush, you're focusing too much on proving, rather than explaining

OpenStudy (anonymous):

Thats the most non-technical way i can think of explaining it

hero (hero):

Kush had it with his very first explanation. He just needs to use numbers and explain each step.

OpenStudy (anonymous):

so finally this is the proof Since... x^a*x^b=x^(a+b) --------------- laws of exponents hence x^a*x^0=x^(a+0)=x^a if x^a*x^0=x^a then, x^0 is a multiplicative identity, and must be equal to 1. EXPLANATION : \[x^a * x^b = x^{(a+b)} \] put x^0 at the place of x^b \[x^a * x^0 \] \[x^a * x^0 = x^{(a+0)} = x^a\] multiplicative identity is the identity which when multplied to any number the answer is the number itself hence here x^0 is the multplicative identity since it is multiplied to x^a and the result is the same x^a Multiplicative identity is 1 but here it is x^0 that implies that x^0 = 1

OpenStudy (anonymous):

i hope this is sufficient for explaining a 12 year average child

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

\[1 = \frac{2^2}{2^2}\] When you are dividing numbers with exponents you subtract the top exponent from the bottom exponent. so \[1 = 2^{2-2}\] so \[1 = 2^0\] This is what kush originally said but i used numbers instead of letters to make the explination clearer

hero (hero):

Good tyler....and kush as well. I think a 12 year old can understand that. Once he does, then you can introduce the more general proof.

OpenStudy (anonymous):

Thank you both - Tyler that helps so much!!!!

OpenStudy (anonymous):

yea u r right good tyler also